OpenStudy (anonymous):

Sheila went to the store and bought 28 apples. She bought twice as many red apples as she bought green, and she bought one-fourth as many yellow as she did red. How many of each type of apple did she buy? A. 16 red; 8 green; 4 yellow B. 12 red; 13 green; 3 yellow C. 14 red; 7 green; 7 yellow D. 18 red; 8 green; 2 yellow

OpenStudy (anonymous):

Only 2 answer choices have twice as many reds as green. Test both of those against the info in the problem to see which one works. Much easier than setting up algebraic equations!

OpenStudy (anonymous):

i dont get you

OpenStudy (anonymous):

It says "She bought twice as many red apples as green apples" Which 2 answer choices match that bit of information?

OpenStudy (anonymous):

well isnt this considered math of any kind

OpenStudy (anonymous):

a and c

OpenStudy (anonymous):

yes... so we know it's a or c Now, it says she bought one fourth as many yellow as she did red...a or c ?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

it would be a

OpenStudy (anonymous):

a) 16 red 4 yellow c) 14 res 7 yellow Is 4 one fourth of 16? Or is 7 one fourth of 14?

OpenStudy (anonymous):

yes...a

OpenStudy (anonymous):

hey thanks!!

OpenStudy (anonymous):

can you help me on a last one!! plz

OpenStudy (anonymous):

no problem....don't do algebra if you don't have to...haha

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

John has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $26.70. Let x represent the number of quarters in the jar. Which equation can be used to find the number of quarters? A. 0.05x + 0.25(150 – x) = 26.70 B. 0.25x + 0.05(150 – x) = 26.70 C. 0.25x + 0.05(x – 150) = 26.70 D. 0.05x + 0.25(x – 150) = 26.70

OpenStudy (anonymous):

Value of Quarters + Value of Nickels = 26.75 0.25x + 0.05(150 - x) = 26.75

OpenStudy (anonymous):

can you explain it to me im trying to understabd because i can never get these when we correct our homework

OpenStudy (anonymous):

x = number of quarters Lets make y = number of nickels OK x + y = 150....it says there are a total of 150 coins, right? AND 0.25x + 0.5y = 26.75...This equations says the value of the quarters plus the value of the nickels is $26.75. Make sense? 2 equations, 2 unknowns...but we can solve the first one easily for y to get y = 150 - x So substitute 150 - x for y in the second equation...and that's the answer.

OpenStudy (anonymous):

I goofed. That 2nd equation should be 0.25x + 0.05y = 26.75 ....not 0.5y